OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4.

OpenStudy (anonymous):

3x+3=-2x+2-4 Try it from there.

OpenStudy (cookiemonster2412):

-1

OpenStudy (anonymous):

yup ^^

OpenStudy (anonymous):

3(x+1) = -2(x-1) - 4 3x + 3 = -2x + 2 - 4 3x + 3 = -2x - 2 3x + 2x = -3 - 2 5x = -5 x = -1